OpenStudy (jkbo):

Please help What is the value of a23? [2 1] [1 2 3] [ a11 a12 a13] [5 3] x [6 7 8] =[ a21 a22 a23] a.23 b.14 c.39 d.31

Directrix (directrix):

First, read the attachment to get an idea about how to multiply two matrices. I am including the link so that you can visit the website to study: http://www.mathwarehouse.com/algebra/matrix/multiply-matrix.php

Directrix (directrix):

@jkbo

Directrix (directrix):

(5*3)+(3*8) = ? Check out the online matrix multiplier if you are allowed to use technology in your work. http://easycalculation.com/matrix/matrix-multiplication.php

OpenStudy (jkbo):

15 + 24 = 39 @Directrix sorry I was in the bathroom

Directrix (directrix):

Alrighty. That is correct. @jkbo

OpenStudy (jkbo):

@Directrix so the answer would be C. 39

Directrix (directrix):

Correct. Multiply the second row of the first matrix by the third column of the second matrix. @jkbo

OpenStudy (jkbo):

@directrix can you please help me What is the value of a 11? [1 3] [5 4] [ a11 a12] [7 2] x [3 2] =[ a21 a22 ] A.26 B14 C.25 D.36

Directrix (directrix):

a 11 is the element that will be in the first row and first column of the product matrix.

Directrix (directrix):

1*5 + 3*3 = ?

OpenStudy (jkbo):

5 + 9 = 14

Directrix (directrix):

That's what I got.